OpenStudy (anonymous):

HELP Find the midpoint of the line segment whose endpoints are (7, 5) and (7, 11).

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

I came up with (7,8)... Correct?

hartnn (hartnn):

yes! good :)

OpenStudy (anonymous):

Yay thanks!!!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

I feel special ^_^

hartnn (hartnn):

you definitely are! ;)